OpenStudy (anonymous):

An architect has a set of blueprints to deliver to a client. He prefers to make his own cylindrical containers. If the diameter of the container is 5 inches and the height is 36 inches, how much cardboard does he need? Round your answer to the nearest hundredth.

OpenStudy (cggurumanjunath):

volume of cylinder =pi*r^2*h

OpenStudy (cggurumanjunath):

r=d/2=5/2=2.5 inch h=36 inch pi=3.14

OpenStudy (anonymous):

I got 282.6

OpenStudy (cggurumanjunath):